OpenStudy (anonymous):

cos(theta)= 1 when theta = n*pi, for n an integer...true or false

jimthompson5910 (jim_thompson5910):

n is an integer, so say n = 0 theta = n*pi would then equal what?

OpenStudy (anonymous):

this was the exact question out of my trig book.

jimthompson5910 (jim_thompson5910):

theta = n*pi = 0*pi = 0

jimthompson5910 (jim_thompson5910):

So when theta = 0, what is cos(theta)

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

if n = 1, then theta = n*pi = 1*pi = pi

jimthompson5910 (jim_thompson5910):

if n = 1, then theta = pi if that's true, what is cos(theta)?

OpenStudy (anonymous):

.1719

jimthompson5910 (jim_thompson5910):

make sure you're in radian mode

OpenStudy (anonymous):

i am...am i missing something?

jimthompson5910 (jim_thompson5910):

cos(pi) is not 0.1719

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

so if theta = pi, then cos(theta) = -1

jimthompson5910 (jim_thompson5910):

go back to the original statement, is it true?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

but we've just shown that cos(theta) = -1

OpenStudy (anonymous):

-1 is still an integer

jimthompson5910 (jim_thompson5910):

but the original statement is cos(theta) = 1

jimthompson5910 (jim_thompson5910):

where theta = n*pi

jimthompson5910 (jim_thompson5910):

n is an integer

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

thank you... im an idiot